OpenStudy (anonymous):

what is the slope of the line? -3x+6y=-6 (show how you get the answer)

OpenStudy (pounce0129):

\[-3x+3x+6y=-6+3y\]

OpenStudy (pounce0129):

\[6y=3x-6\]

OpenStudy (pounce0129):

\[\frac{ 6y }{ 6}=\frac{ 3x }{ 6 }\frac{ -6 }{ 6}\]

OpenStudy (pounce0129):

\[y=1/2x-1\]

OpenStudy (pounce0129):

so the slope is 1/2

OpenStudy (pounce0129):

do you get it @BabetButternubs

OpenStudy (anonymous):

lol yes its just i dont have a caculater that does alpha y= right but thank you sooooo much <3 @pounce0129

OpenStudy (pounce0129):

you could just do it by hand, that's what i did

OpenStudy (anonymous):

im not good with dividing plus im reallly sleepy @pounce0129

OpenStudy (pounce0129):

lol, have a good night :)